OpenStudy (anonymous):

Jack and jill compared their results on their last math test. Jack had 40 correct answers and Jill had 50 correct answers. there were 32 questions that both had correct and 6 they both had wrong. How many questions were on the test?

OpenStudy (anonymous):

Copy this question and past it in mathematics, way more chances for you.

Directrix (directrix):

Work (correct or not) Courtesy of Dorian W at: http://tinyurl.com/bl6m4l6 Consider this argument: (40-32) + (50-32) + 6 + 32 = y (y being the number of questions) (their results must minus the 32 first in order to get the number of questions they got correct, but not together) 8 + 18 + 6 + 32 = y y = 64 questions @dkrall